Write the equation of these lines

(through the point (2,3) with slope 5.)


through the point (-4,3) with slope of three fourths.




Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

a)
y=mx+5
y=5x-6
b)
y=3/4x+6

Answer by Alan3354(69443)

Write the equation of these lines
through the point (2,3) with slope 5
y = mx + b
3 = 5*2 + b
b = -7
y = 5x - 7
-----------
through the point (-4,3) with slope of three fourths
y = mx + b
3 = (3/4)*-4 + b = -3 + b
b = 6
y = 0.75x + 6